Parth (parthkohli):

The area of a square is just double its side. We have:\[\rm x^2 = 2x\]so\[x^2 - 2x = 0\]\[x(x - 2) = 0\]The solutions are 0 and 2, but is it possible to have a square with 0 as its side?

OpenStudy (turingtest):

not really. Some mathematicians have called such objects "degenerate" squares, but you can exclude those kinds of answers when they pertain to physical objects.

OpenStudy (turingtest):

I remember myin's "degenerate circle" question :)

myininaya (myininaya):

That reminds me of a question I asked about circles. Can a "circle" with radius zero still be considered a circle? (x-h)^2+(y-k)^2=r^2 (x-h)^2+(y-k)^2=0 So in other words, can a single point on a graph be seen as a circle with radius 0.

myininaya (myininaya):

lol. We were thinking the same thing.

OpenStudy (turingtest):

Yes, I borrowed Zarkon's answer because I liked it :)

Parth (parthkohli):

Hmm...

OpenStudy (anonymous):

whom are you calling a degenerate?

Parth (parthkohli):

If a point has 0 area, then the point doesn't cover any area... but a point still covers *some* area right?

Parth (parthkohli):

Like infinitesimal, but still, I don't agree with the point that a point has zero area. :p

Parth (parthkohli):

It's not a fair point.

Parth (parthkohli):

Yello.

OpenStudy (jiteshmeghwal9):

i don't gt why \(x^2=2x\) @ParthKohli

OpenStudy (unklerhaukus):

a square has two dimensions

OpenStudy (jiteshmeghwal9):

I have studied something like that \(x^2=x \times x\) \(2x=x+x\)

Parth (parthkohli):

@jiteshmeghwal9 See the top-line of my question.

OpenStudy (turingtest):

a point has no are as it is zero-dimensional, and squares (at least non-degenerate ones) require two dimensions as @UnkleRhaukus said

OpenStudy (turingtest):

no area*

OpenStudy (jiteshmeghwal9):

Ohh ! im so stupid i gt it Okay :)

Parth (parthkohli):

So 0 is not a solution for \(x\)?

OpenStudy (turingtest):

in the purely mathematical sense, I would say "yes it is a solution", but as a "square" in the normal sense of the word, or as a physical object, the answer would be "no".

OpenStudy (unklerhaukus):

if a side length is zero it is not a square

Parth (parthkohli):

Okay, I'm talking about this question. So, yes, 0 is not an answer to the word problem?

OpenStudy (turingtest):

@UnkleRhaukus @Zarkon termed such objects "degenerate". I think this is a sketchy question, they should have included the condition that \(x>0\) to avoid the subtleties in the philosophical mathematical implications of a zero-by-zero square.

OpenStudy (unklerhaukus):

yes zero is not an answer to word problem, however it does solve the equation

Parth (parthkohli):

Math is so idiotic.

Parth (parthkohli):

lol. :)

OpenStudy (unklerhaukus):

maths is a tool,

OpenStudy (turingtest):

True.^ Professional mathematicians throw out useless concepts just because they are such. They also make their own theorems and definitions depending on what they feel is called for in a given situation. For example, many great mathematicians have \(defined\) \(0^0=1\) there is no objective mathematical proof of this, but sometimes it is convenient to do such things. These guys just make it up as they go along basically :P